# Dark photon dark matter from flattened axion potentials

### v1

@@ Abstract
Dark photons can be resonantly produced in the early universe via their coupling to an oscillating axion field. However, this mechanism typically requires large axion-dark photon couplings or some degree of fine-tuning. In this work, we present a new scenario in which efficient dark photon production arises from axion potentials that are shallower than quadratic at large field values. For moderately large initial misalignment angles, the oscillation of the axion field can trigger either efficient dark photon production or strong axion self-resonance via parametric resonance. When self-resonance dominates and disrupts the field’s homogeneity, we show that oscillons—localized, oscillating axion field configurations—naturally form and can sustain continued dark photon production, provided the coupling is $\gtrsim\mathcal{O}(1)$. For dark photon mass up to three orders of magnitude below the axion mass, the produced dark photons can account for a significant fraction of the present-day dark matter. We support this scenario with numerical lattice simulations of a benchmark model. Our results further motivate experimental searches for ultralight dark photon dark matter. The simulation code is publicly available at https://github.com/hongyi18/AxionDarkPhotonSimulator.
@@ 1 Introduction
Unlike scalar fields, vector DM possesses a nontrivial polarization structure, leading to qualitatively different behavior. One particularly interesting context in which the vector nature becomes crucial is the formation of solitons—coherent, localized DM configurations that may form in DM halos [17, 18, 19, 20, 21, 22, 23, 24]. Ground-state vector solitons can be partially or maximally circularly polarized, carrying macroscopic quantum spin [17, 18, 19]. If dark photons couple to ordinary photons, such polarization could leave imprints signatures on astrophysical electromagnetic signals [25]. [^3] This feature is also expected for spin-2 DM [26]. The intrinsic polarization of dark photons also offers a unique experimental signature in both astrophysical and terrestrial searches [27, 28, 29].
Despite their rich phenomenology, the production of ultralight dark photons in the early universe remains a longstanding challenge. A natural possibility is that dark photons are generated via a misalignment mechanism analogous to that proposed for axions. However, reproducing the observed relic abundance through this mechanism typically requires nonminimal couplings to gravity [30], which often violate perturbative unitarity or introduce ghost instabilities [31, 32]. [^4] The vector misalignment mechanism was first discussed in ref.[33], although that work derived the cosmological evolution of a vector field incorrectly. It was later recognized that a nonminimal coupling is required to yield the correct DM relic abundance [30, 31]. Under minimal assumptions about their interactions, dark photons can be produced during inflation as isocurvature fluctuations only if their mass satisfies $m_{X}\gtrsim 10^{-5}\,\mathrm{eV}$ [34, 35, 36]. Alternatively, dark photons may be produced through parametric or tachyonic resonance if they couple to an evolving homogeneous scalar field that induces time-dependent effective mass for the dark photon [37, 38, 39, 40, 41, 42]. These production mechanisms, however, typically require very large couplings [37, 38] or entail a degree of fine-tuning [39, 40, 41]. [^5] Large couplings can be avoided if the scalar field is trapped in a false vacuum until well after the Hubble parameter drops below the scalar mass [41]. Nonetheless, this scenario requires fine-tuning to remain consistent with constraints from the neutron electric dipole moment [43, 44] and may be destabilized by quantum tunneling, raising concerns about its viability. An alternative way to avoid large couplings is to invoke an axion field with a sufficiently large initial velocity to compensate for the small coupling [42]. Dark photons may also be produced via the decay of near-global cosmic strings [45].
@@
Large axion misalignment in flattened potentials gives rise to three key effects. First, the large oscillation amplitude of the axion field around the potential minimum can generate multiple broad instability bands for transverse dark photon modes [60]. Second, the delayed onset of axion oscillations reduces the dilution of dark photons caused by Hubble expansion, making the parametric resonance more effective [41]. Third, flattened potentials naturally facilitate the formation of axion oscillons via self-resonance [61, 62, 63, 64], which are localized nonperturbative field configurations with approximately constant amplitudes and long lifetimes [65, 66, 67, 68, 69, 70]. The efficiency of these effects depends on the coupling strength, the relative amplitude of initial inhomogeneous perturbations, and the initial misalignment angles of the axion field. In scenarios where axion self-resonance dominates over resonant dark photon production with broad instability bands, the broad resonance for the dark photon field will be halted due to the developed inhomogeneities. Instead, oscillons can sustain dark photon production as local sources if the axion-dark photon coupling is not too weak, such that the Bose-Einstein statistics could become effective before the produced dark photons leave the objects [71, 72]. The overall production process is schematically illustrated in figure 2.666The physical picture discussed here can, in principle, also be realized for QCD axions with the potential $V(\phi)=m_{\phi}^{2}f_{\phi}^{2}[1-\cos(\phi/f_{\phi})]$. However, achieving this scenario requires the axion field to be initially positioned near the top of the potential and leads to a fine-tuning problem [47].
@@
In what follows, we first discuss the theoretical and numerical setup and the initial conditions in section 2. In section 3, we analyze scenarios in which dark photons are produced via broad/tachyonic parametric resonance sourced by a homogeneous oscillating scalar field. In section 4, we focus on axion self-resonance and demonstrate that dark photons can be produced from localized oscillons. We summarize our results in section 5. Throughout this work, we adopt the natural units $c=\hbar=1$ and the metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ).
@@ 3.2 Relic abundance of axions and dark photons
[…]
@@
[…]
@@ 4.1 Linear instability analysis
For a (quasi-)periodic background $\bar{\phi}(t)$, axion perturbations $\delta\phi$ can experience self-resonance and grow exponentially. The instability diagram in flat spacetime is shown in figure 6, where several broad resonance bands appear. These bands enable the efficient growth of inhomogeneities and facilitate the formation of oscillons. Achieving strong resonance requires field amplitudes satisfying $\phi_{0}/f_{\phi}\gtrsim\mathcal{O}(5)$.
@@ 5 Conclusions
In this setup, dark photons can account for a substantial fraction of the total DM abundance, amounting to up to $\gtrsim\mathcal{O}(10)\%$, for a wide range of mass ratios, $m_{X}/m_{\phi}\sim\mathcal{O}(10^{-3}$1)1)1 ). Our result provides a viable and less fine-tuned mechanism for generating ultralight dark photon DM, consistent with existing cosmological and astrophysical constraints.
In this work we do not assume any particular origin of the dark photon mass (e.g., Higgs, Stueckelberg, or Proca). [^9] If the vector mass is generated via a Higgs mechanism, it is argued that the formation of vortices can disrupt the viability of dark photons as dark matter [81]. We do not concern about this possibility here. Future work may investigate this scenario assuming a particular dark photon mass origin and the associated observable signatures, such as gravitational wave production, impacts on small-scale structure formation, and potential signals in direct or indirect DM detection experiments. These investigations will offer new insights for probing the dark sector, and provide with a new channel for the production of ultralight vector fields.
@@ Acknowledgments
We would especially like to thank Mustafa A. Amin for useful discussions. We would also like to thank Andrew J. Long and Sudhakantha Girmohanta for helpful comments. HYZ and LV acknowledge support by the National Natural Science Foundation of China (NSFC) through the grant No. 12350610240 “Astrophysical Axion Laboratories”. LV also thanks Istituto Nazionale di Fisica Nucleare (INFN) through the “QGSKY” Iniziativa Specifica project. AC is supported by the National Natural Science Foundation of China (NSFC) through the grant No. 12090064. EDS acknowledges support from the FONDECYT project No. 1251141 (Agencia Nacional de Investigación y Desarrollo, Chile). PA acknowledges support from FONDECYT project No. 1251613. LR is supported by the European Union Funds project 101137080 - Astrocent Plus. This publication is based upon work from the COST Actions “COSMIC WISPers” (CA21106) and “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)” (CA21136), both supported by COST (European Cooperation in Science and Technology).
@@ Appendix B Axion-dark photon coupling
One way to achieve a significant axion-dark photon coupling is to incorporate multiple symmetry-breaking scalar fields, similar to what is done in the clockwork mechanism [37, 82]. Another route to enhancement involves mixing between two scalar fields, where only one directly couples to dark photons. In the case of kinetic mixing with coupling $\epsilon$ or mass mixing through a term $m_{12}^{2}\phi_{1}\phi_{2}$, diagonalization of the kinetic or mass matrices induces an effective coupling of the second scalar field to dark photons. The resulting enhancement factor $j$ is of order $\epsilon$ or $m_{12}^{2}/(m_{2}^{2}-m_{1}^{2})$.

### v2

@@ Abstract
Dark photons can be resonantly produced in the early universe via their coupling to an oscillating axion field. However, this mechanism typically requires large axion–dark photon couplings or some degree of fine-tuning. In this work, we present a new scenario in which efficient dark photon production arises from axion potentials that are shallower than quadratic at large field values. For moderately large initial misalignment angles, the oscillation of the axion field can trigger either efficient dark photon production or strong axion self-resonance via parametric resonance. When self-resonance dominates and disrupts the field’s homogeneity, we show that oscillons—localized, oscillating axion field configurations—naturally form and can sustain continued dark photon production, provided the coupling is $\gtrsim\mathcal{O}(1)$. For dark photon mass up to three orders of magnitude below the axion mass, the produced dark photons can account for a significant fraction of the present-day dark matter. We support this scenario with numerical lattice simulations of a benchmark model. Our results further motivate experimental searches for ultralight dark photon dark matter. The simulation code is publicly available at https://github.com/hongyi18/AxionDarkPhotonSimulator.
@@ 1 Introduction
Unlike scalar fields, vector DM possesses a nontrivial polarization structure, leading to qualitatively different behavior. One particularly interesting context in which the vector nature becomes crucial is the formation of solitons—coherent, localized DM configurations that may form in DM halos [17, 18, 19, 20, 21, 22, 23, 24]. Ground-state vector solitons can be partially or maximally circularly polarized, carrying macroscopic quantum spin [17, 18, 19]. If dark photons couple to ordinary photons, such polarization could leave imprints signatures on astrophysical electromagnetic signals [25]. [^3] This feature is also expected for spin-2 DM [26]. The intrinsic polarization of dark photons also offers a unique experimental signature in both astrophysical and terrestrial searches [27, 28, 29, 30, 31, 32].
Despite their rich phenomenology, the production of ultralight dark photons in the early universe remains a longstanding challenge. A natural possibility is that dark photons are generated via a misalignment mechanism analogous to that proposed for axions. However, reproducing the observed relic abundance through this mechanism typically requires nonminimal couplings to gravity [33], which often violate perturbative unitarity or introduce ghost instabilities [34, 35]. [^4] The vector misalignment mechanism was first discussed in ref.[36], although that work derived the cosmological evolution of a vector field incorrectly. It was later recognized that a nonminimal coupling is required to yield the correct DM relic abundance [33, 34]. Under minimal assumptions about their interactions, dark photons can be produced during inflation as isocurvature fluctuations only if their mass satisfies $m_{X}\gtrsim 10^{-5}\,\mathrm{eV}$ [37, 38, 39]. Alternatively, dark photons may be produced through parametric or tachyonic resonance if they couple to an evolving homogeneous scalar field that induces time-dependent effective mass for the dark photon [40, 41, 42, 43, 44, 45]. These production mechanisms, however, typically require very large couplings [40, 41, 42], entail a degree of fine-tuning [43], or neglect the axion self-resonance which could disrupt the desired features [44, 45]. Dark photons may also be produced via the decay of near-global cosmic strings [46]; however, the overall viability of models in which dark photon DM acquires a Higgs mass is challenged by vortex formation [47, 48, 49].
@@
Large axion misalignment in flattened potentials gives rise to three key effects. First, the large oscillation amplitude of the axion field around the potential minimum can generate multiple broad instability bands for transverse dark photon modes [64]. Second, the delayed onset of axion oscillations reduces the dilution of dark photons caused by Hubble expansion, making the parametric resonance more effective [44]. Third, flattened potentials naturally facilitate the formation of axion oscillons via self-resonance [65, 66, 67, 68], which are localized nonperturbative field configurations with approximately constant amplitudes and long lifetimes [69, 70, 71, 72, 73, 74]. The efficiency of these effects depends on the coupling strength, the relative amplitude of initial inhomogeneous perturbations, and the initial misalignment angles of the axion field. In scenarios where axion self-resonance dominates over resonant dark photon production with broad instability bands, the broad resonance for the dark photon field will be halted due to the developed inhomogeneities. Instead, oscillons can sustain dark photon production as local sources if the axion–dark photon coupling is not too weak, such that the Bose-Einstein statistics could become effective before the produced dark photons leave the objects [75, 76]. The overall production process is schematically illustrated in figure 2.555The physical picture discussed here can, in principle, also be realized for QCD axions with the potential $V(\phi)=m_{\phi}^{2}f_{\phi}^{2}[1-\cos(\phi/f_{\phi})]$. However, achieving this scenario requires the axion field to be initially positioned near the top of the potential and leads to a fine-tuning problem [51].
@@
In what follows, we first discuss the theoretical and numerical setup and the initial conditions in section 2. In section 3, we analyze scenarios in which dark photons are produced via broad/tachyonic parametric resonance sourced by a homogeneous oscillating scalar field. In section 4, we focus on axion self-resonance and demonstrate that dark photons can be produced from localized oscillons. We summarize our results in section 5. Throughout this work, we adopt the natural units $c=\hbar=1$ and the metric signature (,+,+,+)(-,+,+,+).
@@ 3.2 Relic abundance of axions and dark photons
where $T_{\mathrm{osc}}$ is the temperature at $H=H_{\mathrm{osc}}$, and we have taken $g_{*}(T_{\mathrm{osc}})=g_{*S}(T_{\mathrm{osc}})$. This expression provides a parametric estimate of the dark photon relic abundance in scenarios where broad or tachyonic resonance dominates the production. As we will show in section 3.3 from fully nonlinear numerical simulations, the energy density of dark photons typically exceeds that of axions by a factor of $\mathcal{O}(10)$ shortly after the resonance saturates, thus $f_{X}/f_{\phi}\sim\mathcal{O}(10)$. Consequently, dark photons can constitute a dominant component of DM, provided that the mass ratio lies in the range $10^{-3}\lesssim m_{X}/m_{\phi}\lesssim 1$. The lower bound ensures that the dark photon relic density exceeds $10\%$ of the total abundance in axions and dark photons, while the upper bound is required for efficient resonance. In particular, the relic abundances of axions and dark photons become comparable when $m_{X}/m_{\phi}\sim 10^{-2}$.
@@
where $k_{\mathrm{phys}}(z)$ is the physical momentum and $k_{\rm prod}\sim 0.2m_{\phi}$ is the peak momentum at production. Taking the mass ratio $m_{X}/m_{\phi}=0.1$ and assuming that dark photons constitute all DM, we find that $m_{X}\gtrsim 10^{-18}\mathrm{eV}$ would be consistent with the Lyman-$\alpha$ constraint on the free-streaming length $\lambda_{\mathrm{fs}}\lesssim 0.1~\mathrm{Mpc}$ [81, 82]. A more precise treatment of the free-streaming effect would require a full calculation of the matter power spectrum using the full dark photon momentum distribution, which we leave for future work.
@@ 4.1 Linear instability analysis
For a (quasi-)periodic background $\bar{\phi}(t)$, axion perturbations $\delta\phi$ can experience self-resonance and grow exponentially. The instability diagram in flat spacetime is shown in figure 6, where several broad resonance bands appear. These bands enable the efficient growth of inhomogeneities and facilitate the formation of oscillons. [^8] The detailed formation process (including how perturbations backreact on the homogeneoues background and distribute into oscillons) is irrelevant to our disucssions since oscillons are attractor solutions and intermediate processes do not affect their final configurations significantly. For interested readers, see [69, 70, 71]. Achieving strong resonance requires field amplitudes satisfying $\phi_{0}/f_{\phi}\gtrsim\mathcal{O}(5)$.
@@ 5 Discussion and conclusions
In this setup, dark photons can account for a substantial fraction of the total DM abundance, amounting to up to $\gtrsim\mathcal{O}(10)\%$, for a wide range of mass ratios, $m_{X}/m_{\phi}\sim\mathcal{O}(10^{-3}$$1)$. Our result provides a viable and less fine-tuned mechanism for generating ultralight dark photon DM with masses $m_{X}\gtrsim 10^{-18}\mathrm{eV}$ as the dominant DM component, while remaining consistent with existing cosmological and astrophysical constraints.
Our scenario remains valid under different hierarchies of initial fluctuations between the axion and the transverse components of the dark photon field, denoted $\delta\phi$ and $\delta X_{\pm}$. The most natural expectation is $\delta\phi\gg\delta X_{\pm}$ since the axion acquires relatively large curvature fluctuations from its homogeneous background. In this regime, axion self-resonance is highly efficient unless the coupling is sufficiently large, and dark photon production can persist within oscillons. Additional sources of field fluctuations include, for example, inflationary fluctuations (for axions only) and vacuum fluctuations. Depending on the hierarchy, dark photon production may proceed through either channel discussed earlier. See section 2 for more details on initial conditions.
From the point of view of ultraviolet completions, achieving an effective axion-dark photon coupling of order unity in eq. (2.1) can be done naturally in several models. As we discuss in appendix B, a dimensionless parametrization of the coupling in a U(1) dark sector includes an anomaly coefficient $C_{\mathrm{D}}$, a dark fine-structure constant $\alpha_{\mathrm{D}}$, and an enhancement factor $j$ coming from additional dynamics, such as the clockwork mechanism [40, 85, 86] or mixing with another scalar. New states introduced in these ultraviolet completions should align with collider, cosmological or perturbative consistency requirements.
In this work we do not assume any particular origin of the dark photon mass (e.g., Higgs, Stueckelberg, or Proca), since dark photon production in our scenario relies only on the presence of a nonzero mass term for the vector field. However, if the mass arises through a Higgs mechanism, additional effects may become relevant. In particular, the presence of a dark Higgs field opens the possibility of topological defect formation, such as vortices or cosmic strings, which can affect the abundance, phenomenology, and stability of dark photon DM. In fact, vortex formation has been argued to jeopardize the viability of Higgsed dark photon DM [47, 48, 49], making alternative realizations such as the one explored here especially compelling. While our results on axion-induced resonance remain valid in the Higgs case, a full treatment would require incorporating the additional dynamics of the dark Higgs sector. Future work may investigate this scenario and the associated observable signatures, such as gravitational wave production, impacts on small-scale structure formation, and potential signals in direct or indirect DM detection experiments. These investigations will offer new insights for probing the dark sector, and provide with a new channel for the production of ultralight vector fields.
@@ Acknowledgments
We would especially like to thank Mustafa A. Amin for useful discussions. We would also like to thank Andrew J. Long and Sudhakantha Girmohanta for helpful comments. HYZ and LV acknowledge support by the National Natural Science Foundation of China (NSFC) through the grant No. 12350610240 “Astrophysical Axion Laboratories”. LV also thanks Istituto Nazionale di Fisica Nucleare (INFN) through the “QGSKY” Iniziativa Specifica project. AC is supported by the National Natural Science Foundation of China (NSFC) through the grant No. 12090064. EDS acknowledges support from the FONDECYT project No. 1251141 (Agencia Nacional de Investigación y Desarrollo, Chile). PA acknowledges support from FONDECYT project No. 1251613. LR is supported by Teaming for Excellence grant Astrocent Plus (GA: 101137080) funded by the European Union and the grant ‘AstroCeNT: Particle Astrophysics Science and Technology Centre’ carried out within the International Research Agendas programme of the Foundation for Polish Science financed by the European Union under the European Regional Development Fund. This publication is based upon work from the COST Actions “COSMIC WISPers” (CA21106) and “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)” (CA21136), both supported by COST (European Cooperation in Science and Technology).
@@ Appendix B Axion-dark photon coupling
One way to achieve a significant axion–dark photon coupling is to incorporate multiple symmetry-breaking scalar fields, similar to what is done in the clockwork mechanism [40, 85, 86]. Another route to enhancement involves mixing between two scalar fields, where only one directly couples to dark photons. In the case of kinetic mixing with coupling $\epsilon$ or mass mixing through a term $m_{12}^{2}\phi_{1}\phi_{2}$, diagonalization of the kinetic or mass matrices induces an effective coupling of the second scalar field to dark photons. The resulting enhancement factor $j$ is of order $\epsilon$ or $m_{12}^{2}/(m_{2}^{2}-m_{1}^{2})$.